What is the present value of the following uneven cash flow. 150, 350, 750, -85? The interest rate is 9%
Answer #1

Given

CF0 = 150
CF1 = 350
CF2 = 750
CF3 = -85

where, CFn is the cash flow in period n

r = interest rate - 0.09

PV = CF0 + CF1/(1+r) + CF2/(1+r)2 + CF3/(1+r)3

=> PV = 150 + 350/1.09 + 750/1.092 - 85/1.093 = $1036.72